OpenStudy (anonymous):

Is 8 part of the solution set for the inequality 3w ≥27? Why or why not?

OpenStudy (anonymous):

mmm, nope dear , why? Let's try to work it out ^_^, solve for w :\[3w \ge 27 \rightarrow w \ge \frac{27}{3} \rightarrow w \ge 9\] ^_^

OpenStudy (anonymous):

and 8 is neither a factor of 3 nor of 27 either :)